### 12 Mar 2026 v1

@@ Abstract
Submodular maximization constitutes a prominent research topic in combinatorial optimization and theoretical computer science, with extensive applications across diverse domains. While substantial advancements have been achieved in approximation algorithms for submodular maximization, the majority of algorithms yielding high approximation guarantees are randomized. In this work, we investigate deterministic approximation algorithms for maximizing non-monotone submodular functions subject to matroid and knapsack constraints. For the two distinct constraint settings, we propose novel deterministic algorithms grounded in an extended multilinear extension framework. Under matroid constraints, our algorithm achieves an approximation ratio of $(0.385-\varepsilon)$, whereas for knapsack constraints, the proposed algorithm attains an approximation ratio of $(0.367-\varepsilon)$. Both algorithms run in $\mathrm{poly}(n)$ query complexity, where $n$ is the size of the ground set, and improve upon the state-of-the-art deterministic approximation ratios of $(0.367-\varepsilon)$ for matroid constraints and $0.25$ for knapsack constraints.
@@ 1 Introduction
Submodularity, capturing the principle of diminishing returns, naturally arises in many functions of theoretical interest and provides a powerful abstraction for many subset selection problems such as game theory [DBLP:conf/stoc/Dobzinski11, DBLP:conf/soda/DobzinskiS06, DBLP:journals/siamco/DianettiF20], maximal social welfare [DBLP:conf/stoc/FengH0025, DBLP:conf/focs/LiV21, DBLP:journals/siamcomp/KorulaMZ18], influence maximization [DBLP:conf/stoc/MosselR07, DBLP:conf/icml/ChenSZZ21, DBLP:conf/kdd/KempeKT03], facility location [DBLP:journals/dam/AgeevS99, DBLP:journals/orl/Ageev02], data summarization [DBLP:conf/iccv/SimonSS07, DBLP:conf/cikm/SiposSSJ12, DBLP:conf/nips/TschiatschekIWB14], etc. Maximizing submodular functions subject to various constraints, due to $\mathcal{NP}$-hard property, is a fundamental problem that has been extensively studied in both operations research and theoretical computer science, playing a central role in combinatorial optimization and approximation algorithms. Formally, for a ground set $\mathcal{N}$, the submodular function $f:2^{\mathcal{N}}\rightarrow\mathbb{R}$ defined on $\mathcal{N}$ satisfies the following property: for any subsets $A\subseteq B\subseteq\mathcal{N}$ and any element $u\in\mathcal{N}\setminus B$,
$f(A\cup\{u\})-f(A)\geq f(B\cup\{u\})-f(B).$ (1)
@@
The systematic study of submodular maximization dates back to 1978, when Nemhauser et al. [DBLP:journals/mp/NemhauserWF78] abstracted the concept of submodularity from various combinatorial problems and established the 11/e1-1/e approximation ratio for monotone functions via a greedy algorithm. Since then, numerous combinatorial algorithms have been developed, particularly for cardinality constraints [DBLP:journals/mor/NemhauserW78, DBLP:conf/soda/BuchbinderFNS14], knapsack constraints [DBLP:journals/orl/Sviridenko04, DBLP:journals/mor/KulikST13, DBLP:conf/nips/AmanatidisFLLR20, DBLP:journals/algorithmica/FeldmanNS23, DBLP:journals/orl/KulikSS21], matroid constraints [Fisher1978, DBLP:journals/siamcomp/BuchbinderFG23, DBLP:conf/icalp/HenzingerLVZ23], and the unconstrained case [DBLP:journals/siamcomp/BuchbinderFNS15, DBLP:conf/nips/PanJGBJ14]. A significant advancement came with the introduction of the multilinear extension framework by Vondrák [DBLP:conf/stoc/Vondrak08]. This framework extends a discrete submodular function to a continuous multilinear function $F:[0,1]^{\mathcal{N}}\to\mathbb{R}$,
$F_{ML}(\mathbf{x})=\mathbb{E}[f(R(\mathbf{x}))]=\sum_{S\subseteq\mathcal{N}}f(S)\prod_{i\in S}x_{i}\prod_{j\notin S}(1-x_{j}),$
@@
where $R(\mathbf{x})$ is a random subset of $\mathcal{N}$ in which each element $i\in\mathcal{N}$ is included independently with probability $x_{i}$. By applying a Frank-Wolfe-like continuous optimization algorithm, a fractional solution is obtained, which is then converted into a discrete solution via a rounding procedure such as Pipage Rounding [DBLP:journals/siamcomp/CalinescuCPV11] and Swap Rounding [DBLP:conf/focs/ChekuriVZ10]. This approach achieves a 11/e1-1/e approximation ratio for monotone functions under matroid constraints. Crucially, the optimization component of this framework generally requires only that the constraint be a down-closed convex polytope. Given that lossless rounding techniques exist for cardinality and matroid constraints, and that knapsack constraints admit nearly lossless rounding via contention resolution schemes [DBLP:journals/siamcomp/ChekuriVZ14], the multilinear extension framework supports a versatile algorithmic design that is not tied to specific constraint types. Subsequent research has continuously advanced the approximation ratio for non-monotone submodular maximization, from the long-standing $1/e$ [DBLP:conf/stoc/Vondrak08] to $(0.372-\varepsilon)$ [DBLP:conf/focs/EneN16], $(0.385-\varepsilon)$ [DBLP:journals/mor/BuchbinderF19], and $(0.401-\varepsilon)$ [DBLP:conf/stoc/BuchbinderF24]. Closing the remaining gap between these results and the theoretical upper bound of $0.478$ [DBLP:conf/soda/GharanV11] remains a significant open challenge in the field. Computing the value of the multilinear extension exactly typically requires exponentially many function evaluations, and random sampling is regarded as the only polynomial-query approach for estimating the value of the multilinear extension. Additionally, this randomness presents challeng However, the multilinear extension inherently introduces randomness, since computing its value at a given vector ees for applications where deterministic results are crucial, such as in safety-critical or reproducibility-sensitive environments. Randomized algorithms provide guarantees in expectation but not necessarily in the worst case, making them less predictable in certain scenarios. Designing deterministic algorithms for submodular maximization, which returns the same solution on every run and guarantees an approximation ratio in the worst case rather than only in expectation, has emerged as an important and active research direction in recent years. Broadly speaking, deterministic algorithms for submodular maximization follow two main paradigms: (i) designing algorithms that are deterministic by construction, and (ii) derandomizing existing randomized algorithms while preserving their approximation guarantees.
@@
While the Extended Multilinear Extension framework offers a promising avenue for derandomizing submodular maximization algorithms, exploiting this framework in broader settings presents substantial technical challenges. At a high level, the power of EME critically relies on maintaining a distribution whose support size remains constant throughout the algorithm. This structural requirement fundamentally alters the nature of the optimization process: unlike classical multilinear extension methods, the optimization under EME can no longer be carried out via purely continuous procedures. Instead, each update step must carefully combine continuous adjustments with embedded combinatorial subroutines that preserve the bounded support property. Designing such hybrid steps is nontrivial, as standard continuous optimization techniques are no longer applicable. Moreover, the existing derandomization approach for EME is tightly coupled with the exchange properties of matroids, and the resulting algorithmic framework relies heavily on matroid-specific combinatorial structures. As a consequence, these techniques do not readily extend to other algorithms and constraint families. For more general combinatorial constraints, additional difficulties arise: one must design new rounding procedures that are compatible with the EME framework while simultaneously guaranteeing that the support size remains constant. Balancing these requirements poses a central obstacle to extending EME-based derandomization beyond the matroid setting. These challenges naturally raise the following question:
@@
Given a non-negative submodular function $f:2^{\mathcal{N}}\rightarrow\mathbb{R}_{\geq 0}$ and a matroid $\mathcal{M}=(\mathcal{N},\mathcal{I})$, Algorithm 1 is a deterministic algorithm that uses $O_{\varepsilon}(n^{5})$ queries, retu[34] A. Kulik, R. Schwartz, and H. Shachnai. A refined analysis of submodular greedy. Oper. Res. Lett., 49(4):507–514, 2021. [35] A. Kulik, H. Shachnai, and T. Tamir. Approximations for monotone and nonmonotone submodular maximization with knapsack constraints. Math. Oper. Res., 38(4):729–739, 2013. [36] W. Li and J. Vondrák. A constant-factor approximation algorithm for nash social welfare with submodular valuations. In 62nd IEEE Annual Symposium on Foundations of Com- puter Science, FOCS 2021, Denver, CO, USA, February 7-10, 2022, pages 25–36. IEEE, 2021. [37] E. Mossel and S. Roch. On the submodularity of influence in social networks. In D. S. Johnson and U. Feige, editors, Proceedings of the 39th Annual ACM Symposium on Theory of Computing, San Diego, California, USA, June 11-13, 2007, pages 128–134. ACM, 2007. [38] G. L. Nemhauser and L. A. Wolsey. Best algorithms for approximating the maximum of a submodular set function. Math. Oper. Res., 3(3):177–188, 1978. [39] G. L. Nemhauser, L. A. Wolsey, and M. L. Fisher. An analysis of approximations for maximizing submodular set functions - I. Math. Program., 14(1):265–294, 1978. [40] X. Pan, S. Jegelka, J. E. Gonzalez, J. K. Bradley, and M. I. Jordan. Parallel double greedy submodular maximization. In Z. Ghahramani, M. Welling, C. Cortes, N. D. Lawrence, and K. Q. Weinberger, editors, Advances in Neural Information Processing Systems 27: Annual Conference on Neural Information Processing Systems 2014, December 8-13 2014, Montreal, Quebec, Canada, pages 118–126, 2014. [41] I. Simon, N. Snavely, and S. M. Seitz. Scene summarization for online image collections. In IEEE 11th International Conference on Computer Vision, ICCV 2007, Rio de Janeiro, Brazil, October 14-20, 2007, pages 1–8. IEEE Computer Society, 2007. [42] R. Sipos, A. Swaminathan, P. Shivaswamy, and T. Joachims. Temporal corpus summariza- tion using submodular word coverage. In X. Chen, G. Lebanon, H. Wang, and M. J. Zaki, editors, 21st ACM International Conference on Information and Knowledge Management, CIKM’12, Maui, HI, USA, October 29 - November 02, 2012, pages 754–763. ACM, 2012. [43] X. Sun, J. Zhang, S. Zhang, and Z. Zhang. Improved deterministic algorithms for non- monotone submodular maximization. Theor. Comput. Sci., 984:114293, 2024. [44] M. Sviridenko. A note on maximizing a submodular set function subject to a knapsack. Oper. Res. Lett., 32(1):41–43, 2004. [45] S. Tschiatschek, R. K. Iyer, H. Wei, and J. A. Bilmes. Learning mixtures of submodular functions for image collection summarization. In Z. Ghahramani, M. Welling, C. Cortes, N. D. Lawrence, and K. Q. Weinberger, editors, Advances in Neural Information Processing Systems 27: Annual Conference on Neural Information Processing Systems 2014, December 8-13 2014, Montreal, Quebec, Canada, pages 1413–1421, 2014. 35 rns a set $S$ satisfying the matroid constraint $\mathcal{M}=(\mathcal{N},\mathcal{I})$, and achieves $f(S)\geq(0.385-O(\varepsilon))f(\mathrm{OPT})$.
@@
$0.401-\varepsilon$ $poly(n,1/\varepsilon)$ Random [DBLP:conf/stoc/BuchbinderF24]
@@ 2 Preliminary
In this section, we first introduce some basic definitions and essential lemmas adopted in our algorithm design. Without of the ambiguity, we denote $n=|\mathcal{N}|$ and use $\mathrm{OPT}$ to represent the optimal feasible solution. We do not explicitly consider how the submodular function $f$ is computed; instead, we assume access to a value oracle that returns $f(S)$ for any set $S\subseteq\mathcal{N}$ in $O(1)$ time. In addition, for problems under matroid constraints, we assume access to a membership (or independence) oracle that determines whether a given set is independent in $O(1)$ time. The total number of calls to the value oracle and the membership oracle made by an algorithm is referred to as its query complexity.
@@ 4  $S\leftarrow S+v-u$;
Lemma 13 (Theorem 2.4 in [DBLP:journals/corr/abs-2405-05202]).
@@ 3.2 Deterministic Aided Continuous Greedy
The split algorithm serves to extract a forward direction with the largest marginal gain at the current point, under the constraint that the support size remains bounded. This constraint is crucial from a computational perspective: evaluating the extended multilinear extension value incurs a cost that scales with the size of its support. The split algorithm returns a partition of an independent set. As the number of partitions are bounded, the size of support set for the forward direction is also bounded by $l$. Although the theoretical guarantees for Algorithm 3 was proposed in [DBLP:conf/stoc/BuchbinderF25], for completeness, we also restate the proof based on the Lovász extension, which has some subtle differences from [DBLP:conf/stoc/BuchbinderF25].
1Initialize $T_{1}\leftarrow\emptyset,\;T_{2}\leftarrow\emptyset,\;\dots,\;T_{l}\leftarrow\emptyset$;
2 Let $T\leftarrow\bigcup_{i=1}^{l}T_{i}$;
@@ 4  $\mathcal{N}^{\prime}\leftarrow\{u\in\mathcal{N}\setminus T\mid T+u\in\mathcal{I}\}$;
5  $(u,j)\leftarrow\arg\max_{(u,j)\in\mathcal{N}^{\prime}\times[l]}f(u\mid T_{j})$;
@@ 6  $T_{j}\leftarrow T_{j}+u$;
Given a matroid $\mathcal{M}$ on the ground set $\mathcal{N}$, a submodular function $f$, and an integer $l$ as input, Algorithm 3 outputs disjoint sets $(T_{1},T_{2},\dots,T_{l})$ whose union is a basis of $\mathcal{M}$, satisfying for any set $O\subseteq\mathcal{N}$:
$\sum_{j=1}^{l}f(T_{j}\mid\emptyset)\geq\left(1-\frac{1}{l}\right)\cdot f(O)-\frac{1}{l}\sum_{j\in[l]}f(T_{j})$ (4)
@@
Let $T=\bigcup_{j=1}^{l}T_{j}$. Due to dummy elements, we can assume $|T|=|O|=r$. By the exchange property of matroid, there exists a bijection $h:O\rightarrow T$ such that for any $u\in O\setminus T$, $(T-h(u))+u\in\mathcal{I}$. Furthermore, for $u\in O\cap T$, $h(u)=u$. Let $u_{i}\in T\setminus O$ be the $i$-th element added to $T$ in Algorithm 3. Define $T_{j}^{i}=T_{j}\setminus\{h^{-1}(u_{k})\mid k>i\}$, and let $j_{i}$ denote the index of the set to which $u_{i}$ belongs, i.e., $u_{i}\in T_{j_{i}}$. According to the algorithm, we state:
$f(u_{i}\mid T_{j_{i}}^{i-1})\geq\frac{1}{l}\sum_{j=1}^{l}f(h^{-1}(u_{i})\mid T_{j}^{i-1})\geq\frac{1}{l}\sum_{j=1}^{l}f(h^{-1}(u_{i})\mid T_{j})$
@@
$\sum_{j=1}^{l}f(T_{j}\mid\emptyset)$ $=\sum_{i=1}^{r}f(u_{i}\mid T_{j_{i}}^{i-1})\geq\frac{1}{l}\sum_{j=1}^{l}\sum_{i=1}^{r}f(h^{-1}(u_{i})\mid T_{j})$
$\geq\frac{1}{l}\sum_{j=1}^{l}f\left(\bigcup_{i=1}^{r}\{h^{-1}(u_{i})\}\mid T_{j}\right)\quad\text{(by submodularity)}$
$=\frac{1}{l}\sum_{j=1}^{l}\left(f(T_{j}\cup O)-f(T_{j})\right)$
Let $A$ be a random set where for each $1\leq j\leq l$, $\Pr[A=T_{j}\cup O]=\frac{1}{l}$. Then $\frac{1}{l}\sum_{j=1}^{l}f(T_{j}\cup O)=\mathbb{E}[f(A)]$. Let $p_{u}=\Pr[u\in A]$. Since the $T_{j}$ are disjoint, $p_{u}\leq\frac{1}{l}$ for all $u\in\mathcal{N}\setminus O$, and $p_{u}=1$ for all $u\in O$. By Lemma 10,
$\mathbb{E}[f(A)]\geq\hat{f}(p)\geq\int_{\frac{1}{l}}^{1}f(T_{\lambda}(p))d\lambda\geq\left(1-\frac{1}{l}\right)f(O)$
The first inequality is from Lemma 10, the second from the definition of the Lovász extension, and the third follows from the property of $p$: only elements in $O$ have probability $\frac{1}{l}$ of being in the random set. Combining these inequalities yields (4). ∎
Corollary 16 specializes Lemma 15 to the regime $l=1/\varepsilon$ and rewrites the inequality in terms of total marginal gain with respect to the empty set. As we will show later, this form allows us to establish a lower bound on the improvement obtained by the forward direction induced by the split algorithm. Essentially, Corollary 16 provides a quantitative guarantee on the quality of the direction induced by the split algorithm. When the function $f$ is instantiated as the objective evaluated at the current fractional solution, the left-hand side represents the total marginal gain contributed by the components of the forward direction. The inequality then implies that this direction achieves a sufficiently large increase relative to any independent set $O$, and therefore constitutes a sufficiently good forward direction for the continuous greedy process.
@@
In Algorithm 3, for any $O\in\mathcal{I}$, if $l\geq\frac{1}{\varepsilon}$, then:
$\sum_{j=1}^{1/\varepsilon}f(T_{j}\mid\emptyset)\geq(1-2\varepsilon)\cdot f(O)-(1-\varepsilon)\cdot f(\emptyset)$ (5)
@@
$\begin{array}[]{c}\sum_{j=1}^{1/\varepsilon}(f(T_{j})-f(\emptyset))\geq(1-\varepsilon)f(O)-\varepsilon\sum_{j=1}^{1/\varepsilon}f(T_{j})\\ \Updownarrow\\ (1+\varepsilon)\sum_{j=1}^{1/\varepsilon}(f(T_{j})-f(\emptyset))\geq(1-\varepsilon)f(O)-f(\emptyset)\\ \Updownarrow\\ \begin{aligned} \sum_{j=1}^{1/\varepsilon}(f(T_{j})-f(\emptyset))&\geq\frac{1-\varepsilon}{1+\varepsilon}f(O)-\frac{1}{1+\varepsilon}f(\emptyset)\\ &\geq(1-2\varepsilon)f(O)-(1-\varepsilon)f(\emptyset)\end{aligned}\end{array}$
The last inequality uses $\frac{1-\varepsilon}{1+\varepsilon}\geq 1-2\varepsilon$ for $\varepsilon\in(0,1)$ and assumes $f(O)\geq f(\emptyset)$ for a simpler coefficient form. If $f(O)<f(\emptyset)$, the right-hand side is negative. However, dummy elements ensure $f(T_{j})\geq f(\emptyset)$ since $f(T_{j})$ is non-decreasing when elements are added. ∎

### 16 Mar 2026 v2

@@ Abstract
Submodular maximization constitutes a prominent research topic in combinatorial optimization and theoretical computer science, with extensive applications across diverse domains. While substantial advancements have been achieved in approximation algorithms for submodular maximization, the majority of algorithms yielding high approximation guarantees are randomized. In this work, we investigate deterministic approximation algorithms for maximizing non-monotone submodular functions subject to matroid and knapsack constraints. For the two distinct constraint settings, we propose novel deterministic algorithms grounded in an extended multilinear extension framework. For matroid constraints, our algorithm achieves an approximation ratio of $(0.385-\varepsilon)$, while for knapsack constraints, the proposed algorithm attains an approximation ratio of $(0.367-\varepsilon)$. Both algorithms run in $\mathrm{poly}(n)$ value queries, where $n$ is the size of the ground set, and improve upon the state-of-the-art deterministic approximation ratios of $(0.367-\varepsilon)$ for matroid constraints and $0.25$ for knapsack constraints.
@@ 1 Introduction
Submodularity, capturing the principle of diminishing returns, naturally arises in many functions of theoretical interest and provides a powerful abstraction for many subset selection problems such as game theory [DBLP:conf/stoc/Dobzinski11, DBLP:conf/soda/DobzinskiS06, DBLP:journals/siamco/DianettiF20], maximal social welfare [DBLP:conf/stoc/FengH0025, DBLP:conf/focs/LiV21, DBLP:journals/siamcomp/KorulaMZ18], influence maximization [DBLP:conf/stoc/MosselR07, DBLP:conf/icml/ChenSZZ21, DBLP:conf/kdd/KempeKT03], facility location [DBLP:journals/dam/AgeevS99, DBLP:journals/orl/Ageev02], data summarization [DBLP:conf/iccv/SimonSS07, DBLP:conf/cikm/SiposSSJ12, DBLP:conf/nips/TschiatschekIWB14], etc. Maximizing submodular functions subject to various constraints, due to the $\mathcal{NP}$-hard property, is a fundamental problem that has been extensively studied in both operations research and theoretical computer science, playing a central role in combinatorial optimization and approximation algorithms. Formally, for a ground set $\mathcal{N}$, the submodular function $f:2^{\mathcal{N}}\rightarrow\mathbb{R}$ defined on $\mathcal{N}$ satisfies the following property: for any subset $A\subseteq B\subseteq\mathcal{N}$ and any element $u\in\mathcal{N}\setminus B$,
$f(A\cup\{u\})-f(A)\geq f(B\cup\{u\})-f(B).$
@@
The systematic study of submodular maximization dates back to 1978, when Nemhauser et al. [DBLP:journals/mp/NemhauserWF78] abstracted the concept of submodularity from various combinatorial problems and established the 11/e1-1/e approximation ratio for monotone functions via a greedy algorithm. Since then, numerous combinatorial algorithms have been developed, particularly for cardinality constraints [DBLP:journals/iandc/Skowron17, DBLP:journals/mor/NemhauserW78, DBLP:conf/soda/BuchbinderFNS14], knapsack constraints [DBLP:journals/orl/Sviridenko04, DBLP:journals/mor/KulikST13, DBLP:conf/nips/AmanatidisFLLR20, DBLP:journals/algorithmica/FeldmanNS23, DBLP:journals/orl/KulikSS21], matroid constraints [Fisher1978, DBLP:journals/siamcomp/BuchbinderFG23, DBLP:conf/icalp/HenzingerLVZ23], and the unconstrained case [DBLP:journals/siamcomp/BuchbinderFNS15, DBLP:conf/nips/PanJGBJ14]. A significant advancement came with the introduction of the multilinear extension framework by Vondrák [DBLP:conf/stoc/Vondrak08]. This framework extends a discrete submodular function to a continuous multilinear function $F_{ME}:[0,1]^{\mathcal{N}}\to\mathbb{R}$,
$F_{ME}(\mathbf{x})=\mathbb{E}[f(R(\mathbf{x}))]=\sum_{S\subseteq\mathcal{N}}f(S)\prod_{i\in S}x_{i}\prod_{j\notin S}(1-x_{j}),$
@@
where $R(\mathbf{x})$ is a random subset of $\mathcal{N}$ in which each element $i\in\mathcal{N}$ is included independently with probability $x_{i}$. By applying a Frank-Wolfe-like continuous optimization algorithm, a fractional solution is obtained, which is then converted into a discrete solution via a rounding procedure such as Pipage Rounding [DBLP:journals/siamcomp/CalinescuCPV11] and Swap Rounding [DBLP:conf/focs/ChekuriVZ10]. This approach achieves an approximation ratio 11/e1-1/e for monotone functions under matroid constraints. Crucially, the optimization component of this framework generally requires only that the constraint be a down-closed convex polytope. Given that lossless rounding techniques exist for cardinality and matroid constraints, and that knapsack constraints admit nearly lossless rounding via contention resolution schemes [DBLP:journals/siamcomp/ChekuriVZ14], the multilinear extension framework supports a versatile algorithmic design that is not tied to specific constraint types. Subsequent research has continuously advanced the approximation ratio for non-monotone submodular maximization, from the long-standing $1/e(\approx 0.367)$ [DBLP:conf/stoc/Vondrak08] to $(0.372-\varepsilon)$ [DBLP:conf/focs/EneN16], $(0.385-\varepsilon)$ [DBLP:journals/mor/BuchbinderF19], and $(0.401-\varepsilon)$ [DBLP:conf/stoc/BuchbinderF24]. Closing the remaining gap between these results and the theoretical upper bound of $0.478$ [DBLP:conf/soda/GharanV11] remains a significant open challenge in the field. Computing the value of the multilinear extension exactly typically requires exponentially many function evaluations, and random sampling is widely regarded as the only polynomial-query approach for estimating its value. However, such randomness can be undesirable in applications where deterministic outcomes are required, such as safety-critical or reproducibility-sensitive environments. Randomized algorithms provide guarantees in expectation but not necessarily in the worst case, making them less predictable in certain scenarios.
Designing deterministic algorithms for submodular maximization, which returns the same solution on every run and guarantees an approximation ratio in the worst case rather than only in expectation, has emerged as an important and active research direction in recent years. Broadly speaking, deterministic algorithms for submodular maximization follow two main paradigms: (i) designing algorithms that are deterministic by construction, and (ii) derandomizing existing randomized algorithms while preserving their approximation guarantees.
@@
While the Extended Multilinear Extension framework offers a promising avenue for derandomizing submodular maximization algorithms, exploiting this framework in broader settings presents substantial technical challenges. At a high level, the power of EME critically relies on maintaining a distribution whose fractional support value remains constant throughout the algorithm. This structural requirement fundamentally alters the nature of the optimization process: unlike classical multilinear extension methods, the optimization under EME can no longer be carried out via purely continuous procedures. Instead, each update step must carefully combine continuous adjustments with embedded combinatorial subroutines that preserve the bounded support property. Designing such hybrid steps is nontrivial, as standard continuous optimization techniques are no longer applicable. Moreover, the existing derandomization approach for EME is tightly coupled with the exchange properties of matroids, and the resulting algorithmic framework relies heavily on matroid-specific combinatorial structures. As a consequence, these techniques do not readily extend to other algorithms and constraint families. For more general combinatorial constraints, additional difficulties arise: one must design new rounding procedures that are compatible with the EME framework while simultaneously guaranteeing that the support size remains constant. Balancing these requirements poses a central obstacle to extending EME-based derandomization beyond the matroid setting. These challenges naturally raise the following question:
@@
Given a non-negative submodular function $f:2^{\mathcal{N}}\rightarrow\mathbb{R}_{\geq 0}$ and a matroid $\mathcal{M}=(\mathcal{N},\mathcal{I})$, Algorithm 1 is a deterministic algorithm that uses $O_{\varepsilon}(n^{5})$ queries, returns a set $S$ satisfying the matroid constraint $\mathcal{M}=(\mathcal{N},\mathcal{I})$, and achieves $f(S)\geq(0.385-O(\varepsilon))f(\mathrm{OPT})$.
@@
$0.401-\varepsilon$ $\mathrm{poly}(n,1/\varepsilon)$ Random [DBLP:conf/stoc/BuchbinderF24]
@@ 2 Preliminary
In this section, we first introduce some basic definitions and essential lemmas adopted in our algorithm design. Without ambiguity, we denote $n=|\mathcal{N}|$ and use $\mathrm{OPT}$ to represent the optimal feasible solution. We do not explicitly consider how the submodular function $f$ is computed; instead, we assume access to a value oracle that returns $f(S)$ for any set $S\subseteq\mathcal{N}$ in $O(1)$ time. In addition, for problems under matroid constraints, we assume access to a membership (or independence) oracle that determines whether a given set is independent in $O(1)$ time. The total number of calls to the value oracle and the membership oracle made by an algorithm is referred to as its query complexity.
@@ 4  $S\leftarrow S+v-u$;
Lemma 13 (Theorem 2.4 in [NEURIPS2024_c4e40d31]).
@@ 3.2 Deterministic aided Continuous Greedy
The split algorithm serves to extract a forward direction with the largest marginal gain at the current point, under the constraint that the support size remains bounded. This constraint is crucial from a computational perspective: evaluating the extended multilinear extension value incurs a cost that scales with the size of its support. The split algorithm returns a partition of an independent set. As the number of partitions are bounded, the size of support set for the forward direction is also bounded by $\ell$. Although the theoretical guarantees for Algorithm 3 was proposed in [DBLP:conf/stoc/BuchbinderF25], for completeness, we also restate the proof based on the Lovász extension, which has some subtle differences from [DBLP:conf/stoc/BuchbinderF25].
1Initialize $T_{1}\leftarrow\emptyset,\;T_{2}\leftarrow\emptyset,\;\dots,\;T_{\ell}\leftarrow\emptyset$;
2 Let $T\leftarrow\bigcup_{i=1}^{\ell}T_{i}$;
@@ 4  $\mathcal{N}^{\prime}\leftarrow\{u\in\mathcal{N}\setminus T\mid T+u\in\mathcal{I}\}$;
5  $(u,j)\leftarrow\arg\max_{(u,j)\in\mathcal{N}^{\prime}\times[\ell]}f(u\mid T_{j})$;
@@ 6  $T_{j}\leftarrow T_{j}+u$;
Given a matroid $\mathcal{M}$ on the ground set $\mathcal{N}$, a submodular function $f$, and an integer $\ell$ as input, Algorithm 3 outputs disjoint sets $(T_{1},T_{2},\dots,T_{\ell})$ whose union is a basis of $\mathcal{M}$, satisfying for any set $O\subseteq\mathcal{N}$:
$\sum_{j=1}^{\ell}f(T_{j}\mid\emptyset)\geq\left(1-\frac{1}{\ell}\right)\cdot f(O)-\frac{1}{\ell}\sum_{j\in[\ell]}f(T_{j})$ (3)
@@
Let $T=\bigcup_{j=1}^{\ell}T_{j}$. Due to dummy elements, we can assume $|T|=|O|=r$. By the exchange property of matroid, there exists a bijection $h:O\rightarrow T$ such that for any $u\in O\setminus T$, $(T-h(u))+u\in\mathcal{I}$. Furthermore, for $u\in O\cap T$, $h(u)=u$. Let $u_{i}\in T\setminus O$ be the $i$-th element added to $T$ in Algorithm 3. Define $T_{j}^{i}=T_{j}\setminus\{h^{-1}(u_{k})\mid k>i\}$, and let $j_{i}$ denote the index of the set to which $u_{i}$ belongs, i.e., $u_{i}\in T_{j_{i}}$. According to the algorithm, we state:
$f(u_{i}\mid T_{j_{i}}^{i-1})\geq\frac{1}{\ell}\sum_{j=1}^{\ell}f(h^{-1}(u_{i})\mid T_{j}^{i-1})\geq\frac{1}{\ell}\sum_{j=1}^{\ell}f(h^{-1}(u_{i})\mid T_{j})$
@@
$\sum_{j=1}^{\ell}f(T_{j}\mid\emptyset)$ $=\sum_{i=1}^{r}f(u_{i}\mid T_{j_{i}}^{i-1})\geq\frac{1}{\ell}\sum_{j=1}^{\ell}\sum_{i=1}^{r}f(h^{-1}(u_{i})\mid T_{j})$
$\geq\frac{1}{\ell}\sum_{j=1}^{\ell}f\left(\bigcup_{i=1}^{r}\{h^{-1}(u_{i})\}\mid T_{j}\right)\quad\text{(by submodularity)}$
$=\frac{1}{\ell}\sum_{j=1}^{\ell}\left(f(T_{j}\cup O)-f(T_{j})\right)$
Let $A$ be a random set where for each $1\leq j\leq\ell$, $\Pr[A=T_{j}\cup O]=\frac{1}{\ell}$. Then $\frac{1}{\ell}\sum_{j=1}^{\ell}f(T_{j}\cup O)=\mathbb{E}[f(A)]$. Let $p_{u}=\Pr[u\in A]$. Since the $T_{j}$ are disjoint, $p_{u}\leq\frac{1}{\ell}$ for all $u\in\mathcal{N}\setminus O$, and $p_{u}=1$ for all $u\in O$. By Lemma 10,
$\mathbb{E}[f(A)]\geq\hat{f}(p)\geq\int_{\frac{1}{\ell}}^{1}f(T_{\lambda}(p))d\lambda\geq\left(1-\frac{1}{\ell}\right)f(O)$
The first inequality is from Lemma 10, the second from the definition of the Lovász extension, and the third follows from the property of $p$: only elements in $O$ have probability $\frac{1}{\ell}$ of being in the random set. Combining these inequalities yields (3). ∎
Corollary 16 specializes Lemma 15 to the regime $\ell=1/\varepsilon$ and rewrites the inequality in terms of total marginal gain with respect to the empty set. As we will show later, this form allows us to establish a lower bound on the improvement obtained by the forward direction induced by the split algorithm. Essentially, Corollary 16 provides a quantitative guarantee on the quality of the direction induced by the split algorithm. When the function $f$ is instantiated as the objective evaluated at the current fractional solution, the left-hand side represents the total marginal gain contributed by the components of the forward direction. The inequality then implies that this direction achieves a sufficiently large increase relative to any independent set $O$, and therefore constitutes a sufficiently good forward direction for the continuous greedy process.
@@
In Algorithm 3, for any $O\in\mathcal{I}$, if $\ell\geq\frac{1}{\varepsilon}$, then:
$\sum_{j=1}^{\ell}f(T_{j}\mid\emptyset)\geq\left(1-2\varepsilon\right)\cdot f(O)-\left(1-\varepsilon\right)\cdot f(\emptyset)$ (4)
@@
$\begin{array}[]{c}\sum_{j=1}^{\ell}(f(T_{j})-f(\emptyset))\geq\left(1-\frac{1}{\ell}\right)f(O)-\frac{1}{\ell}\sum_{j=1}^{\ell}f(T_{j})\\ \Updownarrow\\ \left(1+\frac{1}{\ell}\right)\sum_{j=1}^{\ell}(f(T_{j})-f(\emptyset))\geq\left(1-\frac{1}{\ell}\right)f(O)-f(\emptyset)\\ \Updownarrow\\ \begin{aligned} \sum_{j=1}^{\ell}(f(T_{j})-f(\emptyset))&\geq\frac{1-\frac{1}{\ell}}{1+\frac{1}{\ell}}f(O)-\frac{1}{1+\frac{1}{\ell}}f(\emptyset)\\ &\geq\left(1-2\varepsilon\right)f(O)-\left(1-\varepsilon\right)f(\emptyset)\end{aligned}\end{array}$
The last inequality uses $\frac{1-\varepsilon}{1+\varepsilon}\geq 1-2\varepsilon+\varepsilon^{2},\frac{1}{1+\varepsilon}\geq 1-\varepsilon-\varepsilon^{2}$ for $\varepsilon\in(0,1)$ and assumes $f(O)\geq f(\emptyset)$ for a simpler coefficient form. If $f(O)<f(\emptyset)$, the right-hand side is negative. However, dummy elements ensure $f(T_{j})\geq f(\emptyset)$ since $f(T_{j})$ is non-decreasing when elements are added. ∎